OpenStudy (questionaskerduh1243):

I NEED HELP IN 5 Or Less MINUTES WILL FAN AND MEDAL PLEASE! how would you use the additive inverse to find 4 - -5

TheSmartOne (thesmartone):

remember a - (-b) = a + b

OpenStudy (questionaskerduh1243):

i dont get it sorry im not that good at math lately

TheSmartOne (thesmartone):

so let's say you're trying to find the distance between a and b The distance would be a - b  because a will be greater than b since it is to the right of b on the number line but what if this b was a negative number? so let's say b is something like -5 and a is 5 we need to find the distance between the two points 5 - (-5) = ?? We can look at a number line again  The distance from -5 to 5 would be the same as adding the distance from -5 to 0 and 0 to 5 -5 to 0 is 5 0 to 5 is 5 5 + 5 = 10 Now let's look at 5 - (-5) again remember I said that a - (-b) = a + b so 5 - (-5) would be the same as 5 + 5 = 10 So do you understand it now? :)